OpenStudy (maleena_1):

Find the magnitude and direction of the vector. Round the length to the nearest tenth and the degree to the nearest unit.

OpenStudy (maleena_1):



OpenStudy (fortytherapper):

Are we finding the hypotenuse, I assume?

OpenStudy (maleena_1):

the magnitude and direction of the vector

OpenStudy (jdoe0001):

magnitude is how "long" it is, thus as aforementioned, it will be the hypotenuse

OpenStudy (jdoe0001):

to find its "direction", is to find the angle it makes, thus, you're given the x,y values of the vector, so, use the inverse tangent function to get the angle :)

OpenStudy (jdoe0001):

\(\begin{array}{llll} <6&,&4>\\ \ x&,&y \end{array}\qquad \begin{array}{llll} ||<6,4>|| magnitude\implies \sqrt{x^2+y^2}\\ direction\implies angle\implies tan^{-1}\left( \frac{y}{x} \right) \end{array}\)

OpenStudy (maleena_1):

magnitude is 7.2 @jdoe0001